OpenStudy (anonymous):

whats the slope of -3x + 2y = 8

OpenStudy (anonymous):

slope of the equation ax+by=c is m=-a/b here a=-3 b=2

OpenStudy (amistre64):

its the difference between x=0, and x=1 :)

OpenStudy (anonymous):

whats the slope of y = -3/2x+4

OpenStudy (anonymous):

slope of the equation in the slope intercept form with slope m and intercept c is y=mx+c compare with this and find the m.

OpenStudy (anonymous):

i dont understand srry

OpenStudy (anonymous):

let me give you example if i have following equation in slope intercept form y=-5x+6 comparing with standard form y=mx+c slope m will be m=-5 can you find it now ?

OpenStudy (anonymous):

yea i got it now